OpenStudy (anonymous):

Security Company A charges $3,915.00 for equipment and installation and then $6.83 per day to monitor the system and respond to alerts. Security Company B charges only $902.00 for the installation of their equipment, but $14.76 per day to monitor the system and respond to trouble. Which system of equations can be used to determine the number of days it takes, x, for the two companies to charge the same amount, y?

OpenStudy (anonymous):

ez two eqns: y=6.83x+3915 y=14.75x+902 y=6.83x+3915 -y=-14.75x-902 0=-7.92x+3,013 x=380.429292929... bit more than a year

OpenStudy (anonymous):

gee thanks, if you are interested in more points i am uploading more